Zinc crystallizes in HCP structure given its crystal parameters a = 2.66A, C = 1 and its atomic weight = 65.37, calculate the density of Zinc
1
Expert's answer
2019-11-29T09:09:34-0500

The packing fraction for zinc is 0.62. In the hexagonal close packed system (HCP) there are 2 atoms per cell. Hence, the density is


"\rho=\frac{M}{V}=\\
\space\\
=\frac{2\cdot0.62\cdot65.37\cdot1.66054\cdot10^{-27}}{(2.66\cdot10^{-10})^3}=7152\text{ kg/m}^3."